i need an age problem that results to 39 years old and 16 years old :) please
Answer by Theo(13342)

jack is 39 years old.
jill is 16.

in how many years will jack be 3 times as old as jill?

solution is as follows:

x = the number of years.

39 + x = 3 * 16 which becomes:
39 + x = 48
subtract 39 from both sides of the equation to get:
x = 48 - 39
solve for x to get x = 9

39 + 9 = 48
3 * 16 = 48
39 + 9 = 3 * 16
jack will be 3 times as old as jill is today in 9 years.